### v1

# From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context

Abstract
Graph Neural Networks (GNNs) have emerged as powerful tools for learning over structured data, including text-attributed graphs, which are common in domains such as citation networks, social platforms, and knowledge graphs. GNNs are not inherently interpretable and thus, many explanation methods have been proposed. However, existing explanation methods often struggle to generate interpretable, fine-grained rationales, especially when node attributes include rich natural language. In this work, we introduce Logic, a lightweight, post-hoc framework that uses large language models (LLMs) to generate faithful and interpretable explanations for GNN predictions. Logic projects GNN node embeddings into the LLM embedding space and constructs hybrid prompts that interleave soft prompts with textual inputs from the graph structure. This enables the LLM to reason about GNN internal representations and produce natural language explanations along with concise explanation subgraphs. Our experiments across four real-world TAG datasets demonstrate that Logic achieves a favorable trade-off between fidelity and sparsity, while significantly improving human-centric metrics such as insightfulness. Logic sets a new direction for LLM-based explainability in graph learning by aligning GNN internals with human reasoning.
From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context
Peyman Baghershahi1thanks: Equal contribution  Gregoire Fournier [^1] footnotemark: 1  Pranav Nyati2  Sourav Medya1 1University of Illinois Chicago, 2Indian Institute of Technology Kharagpur 1{pbaghe2, gfourn2, medya}@uic.edu; 2pranavnyati26@kgpian.iitkgp.ac.in
1 Introduction
Graph neural networks (GNNs) have witnessed rapid adoption across a wide range of critical real-world applications, including healthcare Zitnik et al. (2018, 2019), drug design Xiong et al. (2021); Liu et al. (2022); Sun et al. (2020), recommender systems Chen et al. (2022), and fraud detection Rao et al. (2021). The high-stakes nature of these domains demands that the predictions made by GNNs be not only accurate but also trustworthy. One powerful approach to building trust in deep learning models is to provide meaningful explanations that justify their predictions Shneiderman (2020). Such explanations, which may highlight important substructures in the input Luo et al. (2020); Schlichtkrull et al. (2021); Yuan et al. (2021), or offer counterfactual scenarios Lucic et al. (2022a); Tan et al. (2022a); Chen et al. (2022). However, generating explanations for GNNs is inherently challenging due to the combinatorial nature of graph data and the joint influence of node attributes and edge connections. These complexities make GNN explainability a non-trivial problem and have led to the development of a wide array of explanation techniques Kakkad et al. (2023a); Guo et al. (2023).
1.1 Need for LLM-based GNN explainers
There has been a growing interest in the integration of GNNs and LLMs to enhance GNN task performance on text attribute graphs (TAGs). Essentially, these methods leverage LLMs in different architectural orders to 1) attain rich representations of the input TAGs, task descriptions, and task-specific few-shot examples for the GNN predictors Liu et al. (2024a); Fang et al. (2024), or 2) reason over the TAGs directly Fatemi et al. (2024) or along with GNN-enhanced soft prompts to make predictions Chen et al. (2024a); Tang et al. (2024). However, there is a lack of focus on using LLMs to explain GNNs predictions and clarify the reasoning behind these black-box models.
Traditional GNN explainers aim to identify influential subgraphs or feature subsets that contribute to predictions Ying et al. (2019a); Luo et al. (2020). While effective for graphs with simple node features, these methods struggle in TAGs, where node information is embedded in natural language. Moreover, often these explanations are in the form of subgraphs that are not human-tangible for domain users. LLMs, on the other hand, excel at reasoning over textual content and can generate coherent, human-interpretable rationales. Thus, their integration into GNNs explanation pipelines can improve the interpretability of the GNN models outputs on TAGs.
1.2 Limitations of Existing LLM-based GNN Explainers
The existing frameworks (please see Sec. 2) that use LLMs to build GNN explanations have the following shortcomings: (1) Template dependency and alignment complexity: Aligning the GNN explainers’ output with the LLM acceptable input necessitates rigid templates, handcrafted scores, or extensive training/fine-tuning of the GNN explainer. (2) Missed GNN internals: Existing approaches fail to directly leverage the rich internal representations of the GNN, making the explanations generic or unfaithful to the internal working of the GNNs. (3) Bias from suboptimal explainers: Involving a pre-existing GNN explainer can bias the reasoning and judgment of the LLM. This is critical when the GNN explainers are suboptimal and they signal noisy information to LLM, while the LLM can potentially better infer and reason the pivotal causes of why the GNN has made specific predictions, particularly for TAGs where the textual attributes carry rich information about the graph entities.
1.3 Our Contributions
In this work, we present Logic (Language mOdel-guided GNN Interpretation with Context), a lightweight, post-hoc explanation framework for text-attributed graphs (TAGs) that seamlessly integrates the representational power of GNNs with the reasoning capabilities of LLMs. Our contributions are as follows:
  • Novel Method: We introduce a method that bypasses traditional GNN explainer modules by directly using LLMs as interpreters of GNN behavior. Logic generates fine-grained, natural language explanations without relying on handcrafted templates or perturbation-based saliency masks, reducing bias from external explainers.
  • Soft Prompt Integration: Logic aligns the internal representations of a trained GNN with the token-level embedding space of an LLM via a novel embedding projector. This projection enables GNN features to be treated as soft prompts in a hybrid prompt that blends structural and textual cues, allowing the LLM to reason over the graph’s latent space.
  • Interpretable Explanation Subgraphs: Unlike prior methods that return subgraph masks, Logic produces interpretable subgraphs based on node-level binary support decisions derived from natural language explanations.
  • Training-Agnostic and Plug-and-Play Deployment: Logic does not require fine-tuning the GNN or the LLM, making it easily deployable on pre-trained models without retraining. This makes it adaptable for real-world applications.
2 Related Work
We describe the related studies on the integrated LLM and GNN explainer methods. The explanation capability of LLMs has been applied to graph problems in task-oriented contexts such as explainable recommendation systems Li et al. (2025); Ma et al. (2024). There has been attempts to guide GNNs using the explanations generated by LLMs, for example, in He et al. (2024a), an LLM initially generates node predictions and textual explanations on the reasoning behind the predictions. These explanations then guide a second LLM to produce new node features, which augment the original text features as input for a GNN predictor.
LLM explanation to enhance explainers. LLMs have been utilized to enhance the explanations of pre-existing GNN explainers, for example Zhang et al. (2024) uses LLMs to evaluate explanations of an explainer, its scores guide the weighted gradients used for optimizing the explainer through Bayesian Variation Inference to address learning bias. Similarly, to improve the understandability of explanations, Pan et al. (2024) first trains a pseudo-label generator LLM that takes explanations from a pre-existing GNN explainers and generates pseudo-labels for them. The generator is tuned in an expert iteration procedure with tailored objectives for faithfulness and brevity, and the optimized generator generates pseudo-labels to fine-tune an explainer LLM to build end-to-end models.
LLM explanation to enhance understanding. One advantage of LLMs is that they can generate natural language explanations of GNNs, which are more interpretable for humans. He et al. (2024b) also proposes a method for generating counterfactual graph explanations for GNNs, which uses a counterfactual autoencoder (CA)—composed of a contrastively pretrained text encoder based on LLM-generated text pairs and a graph decoder—to construct counterfactual graph topologies by LLM-generated counterfactual text pairs. To address hallucination, the framework incorporates a dynamic feedback mechanism to refine the generated counterfactual text pairs by prompting the LLM back again first-round generated counterfactuals. Concurrent to our work Cedro and Martens (2025) suggests a narrative (story-like) explanation for a GNN’s predictions. Given an LLM, the dataset description, and a GNN model’s predictions, along with the explanation generated by a GNN explainer in the form of subgraphs and feature importance matrices, they ask the LLM to narrate the relationship between the neighboring nodes, the edges connecting them, and the node’s features but fall short of having the LLM as a method to generate the explanations. Our work aims at addressing these limitations and producing a GNN internal representation-aware LLM explanation. We present additional related work in Appendix A.
3 Problem formulation
We define a graph as $G=(V,E,X)$, where $V$ is the set of nodes, $E=\{(u,v)\mid u,v\in V\}$ is the set of edges, and $X=\{x_{v}\mid v\in V\}$ is the node feature matrix, with $x_{v}\in\mathbb{R}^{d}$ representing the feature vector of node $v$. We denote by $\mathcal{G}$ the set of graphs in a dataset, and by $\mathcal{V}$ the set of all nodes across the graphs in $\mathcal{G}$.
Definition 1 (Text-Attributed Graph (TAG)).
A Text-Attributed Graph (TAG) is defined as a graph
$G=(V,E,T)$
where $V$ is the set of nodes, $E\subseteq V\times V$ is the set of edges (undirected or directed), $T:V\rightarrow\mathcal{D}$ is a function that assigns a textual document or sequence to each node, where $\mathcal{D}$ is a corpus of natural language texts (e.g., sentences, paragraphs).
Each node $v\in V$ is thus associated with a text document $T(v)$, and the graph structure $E$ represents semantic, relational, or structural links between nodes. TAGs enable learning over both the textual content and the topology of the graph. In this work, we focus on the node classification task.
Node Classification. Given a dataset $\mathcal{G}$ and a graph $G=(V,E,X)\in\mathcal{G}$, where every node $v\in V$ has a label $y_{v}\in\mathcal{Y}$ belonging to one of $C$ classes ($\mathcal{Y}=\{Y_{i}\}_{i=1}^{C}$), the node classification task consists in training a GNN $\Phi$ such that $\Phi(x_{v},G)=y_{v}\;\;\forall G\in\mathcal{G},\forall v\in V$.
Our method aims at explaining the node classification predictions of a GNN with local subgraphs, i.e, small subgraphs around the node staged for classification Ying et al. (2019a). More formally:
Definition 2 (Local Factual Explainer).
Given a GNN $\Phi$ and a graph $G=(V,E,T)\in\mathcal{G}$, a local factual explainer $\Psi$ is a mapping from $(\mathcal{V},\mathcal{G})$ to $\mathcal{S}_{\mathcal{V}}(\mathcal{G})$ associated to $\mathcal{V}^{\prime}\subseteq\mathcal{V}$ verifying:
$\Phi(v,\Psi(v,G))=\Phi(v,G)\;\;\forall G\in\mathcal{G},\forall v\in V\cap\mathcal{V}^{\prime}$ (1)
where $\mathcal{S}_{v}(G)\in\mathcal{S}_{\mathcal{V}}(\mathcal{G})$ denotes the set of all subgraphs of $G$ that contain $v$. $\mathcal{V}^{\prime}$ corresponds to the set of nodes over all graphs for which the equation is verified, and in this case, the explanation $\Psi(v,G)$ is called faithful.
Notice that a complete local factual explainer always exists, as $\Psi=\mathrm{Id}$ verifies Equation 1 for every node. However, such explainers are rarely useful in practice, as they do not isolate concise or interpretable subgraphs. We therefore evaluate explanations based on their size and fidelity, as formally defined in Section 5.1.
Although Definition 2 necessitates faithfulness of the explanation graphs and the existing methods that build explanations for GNNs mostly follow this definition, there are two major problems:
  • It does not necessarily generate concise explanations. Generally, a good factual explanation subgraph is a sparse one Kakkad et al. (2023a); Yuan et al. (2023) which has only the most important features of the graph, i.e., nodes, edges, and node features. This property makes the graph more comprehensible for users.
  • More importantly, this problem definition does not account for human-level understanding of the explanations. For example, retrieving a subgraph as a common form of explanations Ying et al. (2019a); Luo et al. (2020) needs additional information for the domain-experts to understand the results and make them useful in practice.
Nevertheless, these two desirable criteria are indeed compatible and aligned with each other as well as the objective of explainable AI (XAI), which attempts to make black-box AI models more interpretable for humans. Therefore, our goal is to design a method to enhance the understandability of the explanations while making them coherent and concise.
Figure 1: Illustration of Logic’s framework. First, the projector is trained to align GNN node embeddings with the LLM’s embedding space. Next, hybrid prompts are constructed by interleaving projected embeddings (as soft prompts) with natural language tokens. These prompts are then fed into the LLM to produce natural language explanations, which are converted into explanation subgraphs.
4 Our Proposed Method: Logic
To explain the predictions of a trained GNN using an LLM, our method projects node embeddings from the GNN latent space into the LLM embedding space, allowing the LLM to reason about the GNN’s predictions. Rather than relying solely on textual prompts, our approach injects GNN embeddings as soft prompts interleaved with natural language tokens. This enables the LLM to generate context-aware explanations grounded in both the model’s internal representations and human-interpretable features. We present our method in three parts: first, the training of the projector; second, the construction of the hybrid prompt; and third, the generation of the explanation subgraph. An overview of our approach is shown in Figure 1.
4.1 Projection from the GNN
In order for the LLM to explain the GNN model, we aim to build a framework where the LLM is able to understand how the GNN sees each node of the computation tree when classifying a particular node. To that effect, we introduce a projector that maps nodes’ GNN embedding into tokens.
Given a GNN $\Phi$ trained on $\mathcal{G}$ for node classification on the node set $\mathcal{V}$, for a given graph $G=(V,E)\in\mathcal{G}$ and node $v\in V\subseteq\mathcal{V}$, we define $f_{\Phi}(v)$ to be the GNN embedding of node $v$ in graph $G$, i.e the feature vector of $v$ before the last prediction layer. The projector aims at both aligning its input with the GNN embeddings and grounding its output in the natural language context of the nodes, hence considering two different levels of embeddings.
Generally speaking, this approach has similarities with other multi-modal alignment frameworks. For instance, in Radford et al. (2021), the authors use contrastive training to align image and text embeddings in a shared space, while in Tsimpoukelli et al. (2021), images are projected directly into a frozen language model’s embedding space to enable coherent language understanding. Similarly, our designed projector learns to map GNN embeddings into soft prompt tokens that serve as input to an LLM, effectively bridging the GNN’s latent space and the LLM’s token-level semantics.
Formally, we define a projector as a function:
$\Pi:\mathbb{R}^{m}\to\mathbb{R}^{k\times h}$
where $m$ is the dimension of the GNN embedding $f_{\Phi}(v)\in\mathbb{R}^{m}$, $k$ is the number of soft-prompt tokens, and $h$ is the hidden dimension of the LLM’s token embeddings. We write $\Pi(f_{\Phi}(v))=Z_{v}\in\mathbb{R}^{k\times h}$. The projector $\Pi$ is trained to optimize two losses:
  1. 1.
    Context alignment loss: This encourages the average soft token representation to align with the LLM embedding of the natural language text associated with the node:
    $\mathcal{L}_{\text{context}}=-\mathbb{E}_{v\in\mathcal{V}}\left[\left\langle\bar{Z}_{v},\;LLM(v)\right\rangle\right]$
    where $\bar{Z}_{v}=\frac{1}{k}\sum_{i=1}^{k}Z_{v}^{(i)}$ is the mean-pooled projector output (normalized), and $LLM(v)\in\mathbb{R}^{k}$ is the normalized embedding of the text associated with node $v$ obtained from the LLM embedding space. The inner product corresponds to cosine similarity.
  2. 2.
    Contrastive loss with GNN embeddings: This encourages soft prompt representations to preserve the similarity structure of the GNN embeddings as follows:
    $\mathcal{L}_{\text{contrast}}=-\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{V}}p_{vu}^{\Phi}\log p_{vu}^{\Pi}$
    where
    $p_{vu}^{\Phi}=\frac{\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(u)\right\rangle/\tau\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle f_{\Phi}(v),f_{\Phi}(w)\right\rangle/\tau\right)}$
    $p_{vu}^{\Pi}=\frac{\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{u}\right\rangle\right)}{\sum_{w\in\mathcal{V}}\exp\left(\left\langle\bar{Z}_{v},\bar{Z}_{w}\right\rangle\right)}$
    and $\tau>0$ is a temperature hyperparameter. Both similarity scores are computed using cosine similarity between normalized vectors.
    The overall training objective becomes:
    $\mathcal{L}=\beta\;\mathcal{L}_{\text{context}}+(1-\beta)\mathcal{L}_{\text{contrast}}$
    where $1\geq\beta\geq 0$ controls the balance between the two losses.
4.2 Prompt Construction
Given a node $u$, the projector outputs a matrix $\Pi(f_{\Phi}(u))=Z_{u}\in\mathbb{R}^{k\times h}$, which we refer to as a soft prompt embedding. This matrix is injected directly into the LLM’s embedding space and interleaved with natural language token embeddings.
For a node $v$ that we wish to explain, we consider its computation tree for GNN $\Phi$, denoted by $\mathcal{T}^{\phi}_{v}$. $\mathcal{T}^{\phi}_{v}$ is a tree with root node $v$ and depth $L$, the number of layers of $\Phi$, composed of all paths of length $L$ starting at $v$, with repetitions, concatenated to the root $v$. We construct the LLM input as a sequence of embeddings consisting of:
  • Standard token embeddings for the system prompt and user instruction.
  • A segment labeled “Target Paper Embedding Representation,” where we insert the soft prompt matrix $Z_{v}$ for the target node, bracketed by textual markers.
  • An enumeration of the nodes in $\mathcal{T}^{\phi}_{v}$, where each node is encoded by interleaved natural language markers and its own soft-prompt embeddings.
  • Final instructions text prompting the model to provide a summary and binary support decision for each computation tree node based on the embeddings.
The LLM sees the embedded GNN features as native tokens, enabling reasoning over both GNN’s representation and natural language inputs. We give an example of the resulting prompt in D.
4.3 Explanation Generation
To produce the explanations, the LLM first predicts whether each node in the computation tree supports or does not support the target node’s classification. This partitions the nodes of the computation tree into three categories: (i) Supporting nodes, (ii) Opposing nodes, (iii) Neutral nodes, not assigned to either category. The explanation subgraph is then formed by including all supporting nodes and a controlled number of neutral nodes. The inclusion of hallucinated nodes is proportional to the number of supporting nodes relative to the total size of the computation tree, as described in our size constraint formulation.
For each node $v$, the LLM generates a natural language explanation $E_{v}$ considering the influence of nodes of $\mathcal{T}^{\Phi}_{v}$ with regards to the classification of node $v$. We define a label attribution function based on the parsing of $E_{v}$:
$\chi:\mathcal{T}^{\phi}_{v}\to\{-1,0,1\}$
such that for each node $u\in\mathcal{T}^{\phi}_{v}$:
  • $\chi(u)=1$ if $u$ is cited in $E_{v}$ as supporting the classification of $v$,
  • $\chi(u)=-1$ if $u$ is cited as not supporting the classification of $v$,
  • $\chi(u)=0$ if $u$ is not mentioned.
We consider the induced partition of $\mathcal{T}^{\phi}_{v}$:
$S^{+}_{v}$ $=\{u\in\mathcal{T}^{\phi}_{v}\mid\chi(u)=1\}$
$S^{-}_{v}$ $=\{u\in\mathcal{T}^{\phi}_{v}\mid\chi(u)=-1\}$
$S^{0}_{v}$ $=\{u\in\mathcal{T}^{\phi}_{v}\mid\chi(u)=0\}$
To address uncertainty and enforce conciseness, we introduce a refinement function $\rho:2^{\mathcal{T}^{\phi}_{v}}\to 2^{\mathcal{T}^{\phi}_{v}}$ such that the final explanation subgraph is:
$S_{v}=\rho(S^{+}_{v},S^{-}_{v},S^{0}_{v})$
[…]
Mitigating Hallucination. LLM-generated explanations may hallucinate, especially when multiple node embeddings are projected within a single prompt. We address this issue through two complementary mechanisms within our framework:
  • Prompt templating with constrained node sets: To reduce hallucinations, we ensure that our prompts contain a fixed set of nodes. We explicitly instruct the LLM to generate explanations only referencing these nodes, limiting the possibility of hallucinating unsupported nodes.
  • Post hoc filtering: After the LLM generates the explanations, any hallucinated nodes are filtered out in a post-processing step. This filtering process ensures that the final explanation remains relevant and grounded in the actual GNN model’s predictions, as detailed in Section 4.3.
These strategies work together to minimize hallucination and enhance the accuracy and reliability of the generated explanations.
Table 1: Performance comparison with baseline models on the node classification task. Higher fidelity and lower size are better. *The explainer returns all-zero masks on Liar, we consider its output to be the central node only.
Cora WikiCS Liar Amazon
Fidelity Size Fidelity Size Fidelity Size Fidelity Size
Node 73.6% 1 77.8% 1 95.0% 1 73.8% 1
Random (50%) 93.5% 18.1 88.9% 19.5 95.0% 19.0 84.4% 7.24
Random (25%) 79.3% 10.0 86.9% 10.0 95.0% 18.13 73.3% 4.03
GNNExplainer 92.0% 18.38 95.4% 9.7 100.0% 12.4 95.5% 5.63
PGExplainer 86.5% 16.2 86.2% 21.2 95.0%* 1.00* 89.0% 5.73
TAGExplainer 60.0% 1.30 80.0% 1.20 95.0%* 1.00* 73.8% 1.52
LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.07
Our method 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 1.30
5 Experiments
Our method, Logic, acts as a light-weight framework to generate explanations of Gnns especially for TAGs. We demonstrate the effectiveness of Logic in various settings. Our code is publicly available at: https://github.com/pbaghershahi/LOGIC.git.
5.1 Evaluation metrics
We use the following metrics in the experiments:
1) Fidelity: Fidelity captures how well an explainable model reproduces or approximates the GNN model logic through its prediction.
Mathematically, the fidelity of the explainer $\Psi$ of GNN $\Phi$ on dataset $\mathcal{G}$ over $\mathcal{V}$ nodes is:
$Fid=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in V}\mathds{1}(\Phi(v,\Psi(v,G))=\Phi(v,G))$
2) Explanation size: The explanation size measures the complexity or compactness of the explanation subgraph produced by the explainer $\Psi$. Specifically, it captures the number of nodes in the explanation for a given prediction. Smaller explanation sizes are generally preferred, as they are easier to interpret. The explanation size is:
$Size=\frac{1}{|\mathcal{V}|}\sum_{G\in\mathcal{G},v\in\mathcal{V}}|\Psi(v,G)|$
5.2 Experimental setup
Datasets. We use real-world TAG datasets to evaluate the performance of our method: Cora Sen et al. (2008) - a co-citation network of research papers, WikiCS Mernyei and Cangea (2022) - a network of related Wikipedia articles, the Amazon-Product Feng et al. (2024) dataset - an E-commerce product co-purchase network, and the Liar Wang (2017) dataset - a fake news detection knowledge-graph. Additional details about the processing and relevance of the datasets are provided in Appendix B.
Models. We consider a GCN Kipf and Welling (2017a) as the base GNN model for the node classification task for all datasets. We train a model consisting of three convolutional layers, with the last one being the classification layer. The training/validation/testing split is $60\%/20\%/20\%$ for all the datasets. We give additional information about the GCN training in Appendix C. We use Meta’s Llama-3.1-8B-Instruct from the HuggingFace library as the frozen LLM in our experiments. We run all experiments on NVIDIA-GeForce-RTX-3090 GPUs, 24GB RAM.
Figure 2: A visualization of the explanations on the Amazon dataset. Left shows the explanations subgraph generated by GNNExplainer. Right shows the explanation generated by Logic. The node marked by a bold circle (product #7) is the node of interest for the explanation. Nodes with the same color are from the same category. Although both models have similar quantitative performance (in terms of size and fidelity), the LLM-generated explanations are textually clarified, making them more human-understandable.
Baselines. To compare our method’s performance with existing approaches, we consider the following five baselines: (1) Node: This baseline uses only the target node itself as the explanation subgraph. This represents the minimal possible context and serves as a lower bound on explanation size. (2) Random: We introduce a randomized baseline that selects a subgraph around each target node as its explanation subgraph. Specifically, for each node, we sample a subgraph of half the size of its computation tree. The results are averaged over 5 samples. (3-5) Graph Explainers: We compare Logic with GNNExplainer Ying et al. (2019a), PGExplainer Luo et al. (2020), subgraph-based explainers for node classification tasks, which do not involve any LLM-based component.
We add the recent LLM-based graph explanation method, TAGExplainer Pan et al. (2024). To the best of our knowledge, TAGExplainer is the only existing LLM-based explanation approach specifically designed for TAGs.
5.3 Quantitative Retults
Table 1 presents the results. One major observation is that on all the datasets, the Node baseline (target node as the singleton explanation subgraph) is considerably powerful, denoting how informative the textual attributes of each node are to guide the GNN’s predictions. Assessing based on fidelity, GNNExplainer, achieves the best performance over most of the datasets, except for Cora, on which Logic achieves better results. However, as discussed before, fidelity alone does not reflect the superiority of a method. Essentially, conventional GNNs (GCN in our case) score a sample, i.e., a node, based on the combination of its own information and the aggregated information of its neighbors, meaning that the context surrounding a node highly impacts its prediction score. Therefore, a GNN explainer that explains a target node with its whole induced subgraph achieves 100% fidelity. Meanwhile, the majority of nodes in the same neighborhood of graphs with a high homophily ratio belong to the same category. Thus, selecting a comparably large number of neighbors of a target node on these types of graphs can most probably guarantee the same representation for the target node as if the whole induced subgraph is selected.
Logic achieves a strong balance between fidelity and explanation size across all four datasets. Notably, Logic matches or exceeds the fidelity of traditional explainers like GNNExplainer, outperforming it on Cora, while producing substantially smaller explanation subgraphs, particularly on Liar and Amazon. This result is especially significant for TAGs, where textual node attributes carry rich semantic information; Logic effectively distills these semantics into compact, interpretable explanations without compromising on alignment with the original GNN predictions. In contrast, baseline methods such as Random or Node either inflate the explanation size or overly restrict the context. These results validate our design choice of using LLMs guided by projected GNN embeddings to construct context-aware yet concise explanation subgraphs, offering a superior trade-off between interpretability and predictive fidelity.
We present additional experiments for different GNN architectures and LLMs in Appendix E.
Table 2: Expert qualitative evaluation of the explanations from Logic and GNNexplainers on Cora. $M_{1}$ denotes the mean scores for GNNexplainer and $M_{2}$ the mean scores for Logic.
Dimension $M_{1}\uparrow$ $M_{2}\uparrow$ $\Delta(M_{2},M_{1}$)
Understandability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
Trustworthiness $2.7_{\scriptstyle\pm 0.8}$ $3.1_{\scriptstyle\pm 0.8}$ $0.4_{\scriptstyle\pm 0.8}$
Insightfulness $3.3_{\scriptstyle\pm 1.1}$ $3.1_{\scriptstyle\pm 1.1}$ $0.2_{\scriptstyle\pm 1.1}$
Satisfaction $2.2_{\scriptstyle\pm 0.9}$ $2.7_{\scriptstyle\pm 0.9}$ $0.5_{\scriptstyle\pm 0.9}$
Confidence $0_{\scriptstyle\pm 0.9} 3.3$$3_{\scriptstyle\pm 0.9}$ $0.3_{\scriptstyle\pm 0.9}$
Convincingness $3.3_{\scriptstyle\pm 1.2}$ $3.7_{\scriptstyle\pm 1.2}$ $0.4_{\scriptstyle\pm 1.2}$
Communicability $2.9_{\scriptstyle\pm 0.8}$ $3.2_{\scriptstyle\pm 0.8}$ $0.3_{\scriptstyle\pm 0.8}$
Usability $3.3_{\scriptstyle\pm 0.4}$ $3.5_{\scriptstyle\pm 0.4}$ $0.2_{\scriptstyle\pm 0.4}$
5.4 Qualitative Evaluation
Results on Cora. We follow the qualitative evaluation framework provided in Cedro and Martens (2025), and refer to Appendix F for a full description of the evaluation dimensions. An expert reviewed the explanations from GNNexplainer and from our method over 10 node classification explanation tasks, and the evaluation framework is developed in Appendix G. The results are presented in Table 5.3. We observe that our method appeared more trustworthy and convincing to our expert, which may be explained by the semantic analysis abilities of the LLM of our graph explaining method. Concurrently, the explanations were also less insightful, as this may have limited the LLM in its exploration of topics and application not directly limited to the subject of the research paper.
Visualization on Amazon dataset. Figure 2 illustrates a representative example of Logic output on the Amazon dataset to emphasize the narrative quality and interpretability of the generated explanation. In this instance, the target (bold) node corresponds to a product (ID 7), predicted to belong to the “Sports & Outdoors” category. Logic analyzes the local neighborhood of the target node and translates the latent GNN-based representations into natural language explanations by leveraging the LLM’s language generation capabilities, interleaved with soft GNN prompts.
Logic produces node-wise rationales from the computation tree of the target node, explicitly stating for each neighbor a brief semantic summary along with a binary support judgment (YES/NO) regarding its influence on the target’s classification.
Table 3: Performance comparison for the ablation study. LLM denotes prompting the LLM with words only and no projected embeddings. LLM+Pr is LLM with a projector. LLM with projector and post-processing (denoted by LLM+Pr+Po) is our method.
Cora WikiCS Liar Amazon
Fidelity Size Fidelity Size Fidelity Size Fidelity Size
LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.2
LLM+Pr 92.2 % 16.0 80.1% 3.7 95.0% 10.1 86.5% 1.7
LLM+Pr+Po 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 4.9
5.5 Ablation Study
To assess the effectiveness of the different components of our approach, we measure the performance of our method with and without the projector and the post-processing. The results are in Table 3. We observe that both fidelity and explanation size drop across all datasets when we replace the words in the prompt with the projected GNN embeddings. This might be due to the LLM having to extract knowledge from more abstract representations, making it more selective in its reasoning. We also observe that adding the post-processing step increases fidelity, at a reduced cost in size compared to the method without a projector, on the WikiCS dataset. From this analysis, we suggest that our method offers a good compromise: achieving high fidelity (slightly lower than directly prompting the LLM with node-associated text) while maintaining a compact explanation size (slightly larger than projected GNN embeddings without post-processing).
6 Conclusions
In this work, we propose Logic, a lightweight, post-hoc explanation framework that uses LLMs to generate faithful and human-interpretable explanations for GNN predictions on text-attributed graphs. By projecting GNN embeddings into the LLM space, Logic enables context-aware, natural language rationales without relying on traditional GNN explainer modules that often need training. Our approach produces concise explanation subgraphs with high fidelity and semantic clarity, outperforming prior methods. Logic demonstrates that LLMs can serve as effective interpreters for GNNs and offers a scalable and intuitive path toward explainability in graph-based applications.
7 Ethical considerations
In this work, we have built methods for generating explanations for graph neural network predictions. We do not foresee any ethical issues from our study.
8 Limitations
We present some limitations of our work:
  • Our method projects GNN embeddings into the LLM embedding space and interleaves them with text to form hybrid prompts. However, batching multiple nodes leads to degraded performance and hallucinations, likely due to interference between soft prompts.
  • While Logic achieves high empirical fidelity, it does not provide formal guarantees of faithfulness to the GNN’s internal decision logic, as the explanation relies on the LLM’s interpretation of projected embeddings rather than directly tracing GNN computation paths.
  • Although our method incorporated a projector to align the embeddings of the GNN with the LLM space of token embeddings, the objectives we introduce for training the projector are not directly tailored for the task of explanation. More optimal objectives that adhere to the explanation task could bring substantial benefits.
Appendix A Additional Related Work
Here we provide additional related works that are on the integrated methods of LLMs and GNNs. The integration of LLMs with GNNs aims to leverage the advanced reasoning abilities of LLMs and the extensive background knowledge they acquire, alongside the GNNs’ proven capacity to enhance the exploitation of graph structures via message-passing mechanisms Kipf and Welling (2017b); Hamilton et al. (2017). The approaches for integrating LLMs and GNNs primarily vary in the sequence of component application Ren et al. (2024); Chen et al. (2024b).
LLMs as prefix and postfix. First, there are methods that utilize LLMs at the prefix modules of their frameworks for task and feature representation Liu et al. (2024a); Fang et al. (2024); He et al. (2024a); Yang et al. (2025). This architecture provides greater flexibility, as various tasks can be specified in natural language with text-based graph representations, and LLM modules can remain fixed. On the other hand, the second class of methods uses GNNs as the prefix Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024), which is better suited for capturing a graph’s topology and its intricate relationships and uses LLMs for prediction. This approach also offers flexibility, as the LLM can output embeddings, prediction scores, or textual sequences to address various downstream problems specifically for generation tasks Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024).
Independent LLMs and joint LLM-GNNs. Additionally, there is a body of work that employs LLMs independently to tackle graph-related tasks Cheng et al. (2024); Fatemi et al. (2024); Li et al. (2024). Approaches following this paradigm benefit from the generalization capabilities of LLMs at both input and output levels, and can potentially avoid training learnable parameters by fixing the LLM component and relying on input prompts. Finally, there is a group of works which employ Gnns and LLMs jointly through soft prompting, which necessitates alignments between their embeddings while enhancing LLMs reasoning by the Gnn captured graph structural features Liu et al. (2024b); Wang et al. (2024); Tian et al. (2023).
Explainers for GNNs. Several algorithms have been proposed in the literature to enhance the interpretability of GNNs. The majority of existing explainers concentrate on instance-level explanations. Instance-level (or local) explainers operate on individual graphs, identifying influential components—such as subgraphs—that most strongly impact the model’s prediction Ying et al. (2019b); Luo et al. (2020); Shan et al. (2021); Huang et al. (2022); Yuan et al. (2022); Lucic et al. (2022b); Tan et al. (2022b); Lin et al. (2021); Bajaj et al. (2021); Abrate and Bonchi (2021); Chhablani et al. (2024); Wellawatte et al. (2022); Verma et al. (2024); Lu et al. (2024). However, this localized perspective constrains their ability to uncover global patterns leveraged by GNNs across multiple graphs, as well as how these patterns integrate into a unified decision-making process. Research on global GNN explainers remains relatively limited Yuan et al. (2020); Huang et al. (2023); Xuanyuan et al. (2023); Azzolin et al. (2023); Armgaan et al. (2024); Fournier and Medya (2025). For instance, XGNN Yuan et al. (2020) and GNNInterpreter Wang and Shen (2023) adopt generative modeling approaches, producing graphs that most strongly correspond to a given class label. For more details, please refer to this survey on GNN explainers Kakkad et al. (2023b).
Table 4: Statistics of the graph datasets.
Dataset # Nodes # Edges # Categories # Features
Cora 2,708 5,429 7 2000
WikiCS 11701 148555 10 300
Amazon 1000 1397 47 2000
Liar 13293 32443 7 387
Appendix B Datasets
B.1 Datasets Details
We describe in detail the datasets that we used in our evaluation below. The basic statistics about these datasets are provided in Table 4.
  1. 1.
    Cora: It is a citation network, in which nodes represent computer science research papers, and each edge between two nodes represents a research paper citing another. The nodes are classified into one of seven categories. Though a citation network is a directed network, the dataset is widely used as an undirected network in the message-passing based graph machine learning, especially for the node classification task, in which the task is to predict each node’s category based on its own text features and text features of its neighbors. Each node’s feature is a 2000-dimensional bag-of-words representation of the keywords in the paper that it corresponds to.
  2. 2.
    WikiCS: It is a text-attributed graph dataset, derived from the Wikipedia platform, widely used for node classification tasks. The nodes correspond to Wikipedia page descriptions of different computer science topics, and edges between nodes represent hyperlinks from one article to another. Each node in the dataset belongs to one of 10 categories. The node classification task is to correctly predict a node’s label, and the edges are considered to be undirected. Each node’s feature vector is a GloVe embedding of size 300, computed based on the text associated with the nodes.
  3. 3.
    Liar: It is a fake-news detection dataset that is often represented as a knowledge graph, with nodes corresponding to statements, speakers, and topics, and edges encoding typed relations such as spoken_by and about.To adapt Liar   into a homogeneous graph suitable for standard GNN pipelines, we merge the three node types—statements, speakers, and topics—into a single unified node set. Each node is embedded with a 384-dimensional BERT representation and augmented with a 3-dimensional one-hot vector indicating its type. Edges representing the original heterogeneous relations (spoken_by and about) are converted to undirected edges and unified into a single edge type in the homogeneous graph. Statement nodes retain their ground-truth labels for fake news classification (ranging from pants-fire to true), while speaker and topic nodes are assigned a dummy label. We also preserve a node-level string attribute—containing the raw statement, speaker name, or topic—for use in explanation and visualization.
  4. 4.
    Amazon-Product: It is a network with nodes representing different kinds of products on Amazon and edges connecting co-purchased products. The nodes belong to one of 47 different categories of products. We use only a subset of the Amazon-Product   dataset, consisting of the first 1000 products and their co-purchase edges, as the entire dataset is very large, and call it Amazon   in our work. The node classification task is to predict a product’s category based on its own as well as its neighbours’ features.
B.2 Relevance of datasets
Table 5: Dataset Descriptions and Relevance to Our Work
Dataset Description & Relevance Why It Fits Our Work
CORA Citation network of CS papers with bag-of-words features. Standard benchmark for node classification on text-rich graphs.
WIKICS Wikipedia CS articles linked by hyperlinks, with GloVe embeddings. Large, semantic-rich graph ideal for explanation evaluation.
LIAR Fake news detection graph combining statements, speakers, topics. Challenging, heterogeneous graph testing method adaptability.
AMAZON Product co-purchase network with 47 categories. Real-world, diverse e-commerce graph to test scalability.
The datasets used in our work are selected based on their relevance to graph explainability tasks, allowing us to assess the effectiveness and robustness of our methods across a diverse set of graph structures and features. The chosen datasets span a variety of domains, including citation networks, text-attributed graphs, fake-news detection, and e-commerce, which reflect both real-world complexity and common challenges faced in explainability. The characteristics of each dataset make them suitable for evaluating the performance of our explanation method. Table 5 summarizes the key features of the datasets and explains their relevance.
Appendix C Additional Implementation Details
We discuss additional details related to the training of the base GNN.
C.1 GNN training
For all datasets, we implement a GCN with 3 convolutional layers, with the input and output channel dimensions chosen according to each dataset’s feature dimension and number of classes, respectively. For the hidden dimension, we choose either $64$ or $512$, depending on the nodes’ feature dimension. The model is trained with the Adam optimizer Kingma and Ba (2014) and a learning rate of 0.001 for 400 epochs.
C.2 GNN accuracy
We give the GNNs accuracy for the different folds in Table 6.
Table 6: Accuracy of GCN on the node classification task for the Cora, Wiki CS, Liar and Amazon products datasets.
Dataset Training Validation Testing
Cora 0.88 0.81 0.84
WikiCS 0.89 0.82 0.81
Amazon 0.87 0.73 0.76
Liar 0.42 0.30 0.29
Appendix D Prompt and generated responses examples
In this section we present a detailed prompt construction in D.1 and a complete example of such built prompt with its input in D.2.
D.1 Example of Full Prompt Construction for Logic
We have provided an example of the full prompt used in Logic below. The boldface texts denote where the embeddings are injected into the prompt. An example response is also illustrated in Figure 3 in the paper.
Prompt:
You are an expert research assistant. Your goal is to assess whether the classification of a target product is supported by its co-purchase neighborhood by analyzing Amazon product reviews and its co-purchase neighborhood to understand why it has been classified under a specific category. The categories to distinguish from are Home & Kitchen, Health & Personal Care, Beauty, etc.
Target Product ID: 0
Predicted Category: Clothing, Shoes & Jewelry
Target Product Embedding Representation:
[…]
Neighboring Products in the Co-purchase Network:
Each product below is described by keywords.
  • Product 129:
[…]
  • Product 130:
[…]
Instructions:
You are given a target product and a list of neighboring products, each described by keywords.
For each neighboring product:
  1. 1.
    Write one sentence summarizing the main topics or ideas captured in its keywords.
  2. 2.
    Clearly state whether this product supports the classification of the target product into category ‘Clothing, Shoes & Jewelry’.
Use the following format for each neighbor:
Product {ID}:
Summary: One sentence summary of the product’s keywords.
Support: YES or NO — Does this product support the classification into ‘Clothing, Shoes & Jewelry’?
Base your reasoning only on the keywords and proximity to the target product.
Start your analysis below:
D.2 Sample of the prompt and output for the Amazon-Product dataset
We provide a sample of the prompt and a sum up of the output we obtain (without the supporting sentence for each product) in Figure 3.
Figure 3: An illustration of the prompting template and LLM explanation of Logic for the Amazon dataset. The left block shows the input prompt to the LLM for a sample target node (target node id 3). The right block denotes the LLM response explaining the prediction of the GNN for the target node.
Appendix E Additional Experiments
We extend Table 1 of the paper with additional experiments for Amazon with TAGExplainer Pan et al. (2024), PGExplainer Luo et al. (2020) and with the GIN Xu et al. and GAT Veličković et al. (2017) architectures. The results are reported in Table 7.
Table 7: Performance comparison on the Products dataset using different GNN architectures (using "Llama 3.1 8B Instruct"). Higher fidelity and lower size are better.
GCN GAT GIN
Fidelity Size Fidelity Size Fidelity Size
Node 73.8% 1.00 73.2% 1.00 39.6% 1.00
Random 84.4% 7.24 84.2% 7.28 90.2% 7.25
GNNExplainer 95.5% 5.63 94.5% 5.615 88.0% 5.65
PGExplainer 87.2% 5.46 88.3% 5.76 87.2% 5.46
TAGExplainer 73.8% 1.37 71.8% 1.43 42.3% 1.27
LLM 86.5% 5.07 91.0% 5.00 86.5% 5.18
Our method 87.0% 1.3 81.5% 1.45 60.0% 1.52
Our method produces the most compact explanations (close to size 1) across all architectures, achieving a strong balance for the GCN with 87.0% fidelity at just 1.3 in size. However, while explanation size remains minimal for the GAT and GIN models, fidelity drops more noticeably, particularly with GIN (60.0%), suggesting that our projection-based mechanism may be less effective when the model relies heavily on structural information or deep aggregation. In contrast, GNNExplainer maintains the highest fidelity across all architectures but at the cost of much larger explanations, while PGExplainer and TAGExplainer underperform in fidelity despite moderate sizes. The LLM baseline shows robust fidelity, especially on GAT, but its explanations are substantially less concise. These results suggest that our method offers an excellent fidelity/interpretability trade-off on GCN, highlighting the need for architecture-aware extensions when applied to more structural-based models.
Table 8: Results on Products dataset using different pretrained LLMs (for pretrained GCN). Higher fidelity and lower size are better.
Llama 3.1 8B Instruct Mistral 7B Instruct v0.2 GPT-Neo 2.7B Phi-3 Mini 4k Instruct Pythia 2.8B
Fidelity Size Fidelity Size Fidelity Size Fidelity Size Fidelity Size
LLM 86.5% 5.07 88.0% 5.14 84.0% 2.33 88.0% 5.17 84.0% 2.34
Our method 87.0% 1.30 83.0% 1.63 80.0% 1.00 80.0% 1.00 80.0% 1.00
The baseline LLM method, which directly leverages the language model without projecting GNN embeddings, generally achieves higher fidelity scores but produces substantially larger explanations, two to three times larger. In contrast, our method consistently generates more compact explanations—typically around size 1, with a moderate decrease in fidelity of roughly 4–7 percentage points depending on the model. Notably, Llama 3.1 and Mistral backbones show strong fidelity for both methods, whereas smaller models like GPT-Neo, Phi-3 Mini, and Pythia exhibit lower fidelity overall. These results highlight our method’s effectiveness at generating precise explanations while maintaining reasonable alignment with the GNN’s predictions and demonstrate its compatibility with different LLM models.
Appendix F Definition of the qualitative metrics
We adapt the different dimensions of the framework in Cedro and Martens (2025) to the context of Cora to clarify their definitions, and provide the definition as given to the expert.
  1. 1.
    Understandability. Explanations should clearly convey why a given paper was assigned to a particular research topic.
  2. 2.
    Trustworthiness. Explanations should help users assess whether the model’s classification of a paper can be trusted.
  3. 3.
    Insightfulness. Explanations should reveal insights about the applications or connections that might play a role in the classification.
  4. 4.
    Satisfaction. Explanations should feel complete and meaningful to users evaluating the model’s behavior.
  5. 5.
    Confidence. Explanations should help users gain confidence in the correctness of the classification.
  6. 6.
    Convincingness. Explanations should be persuasive in justifying the model’s decision for a given paper.
  7. 7.
    Communicability. Explanations should be expressed in a way that aligns with the user’s background knowledge and expectations.
  8. 8.
    Usability. Explanations should support practical tasks such as interpreting predictions, or improving model performance.
Appendix G Experimental Setup for the human evaluation of M1 and M2 Scores
We clarify the experimental setup used to obtain the M1 and M2 scores reported in Table 2.
A) Data Preparation
We randomly selected 10 articles from the Cora dataset, on which a GNN had been trained. For each article, we generated explanations using both LOGIC and GNNExplainer. Each expert was provided with:
  • The original article text and the GNN’s predicted label.
  • Two explanation subgraphs (one per method), anonymized and presented in random order known only to the experimenters.
  • For each subgraph, we listed the texts of the nodes it contains, sorted by increasing node index.
B) Evaluation Procedure
  • Before the evaluation, we provided experts with definitions of the evaluation metrics (as described in Appendix F) and asked them to consider how to apply each metric to content from the Cora dataset.
  • During the evaluation, each expert received the materials described above and independently rated each explanation on a scale from 1 to 5 for each metric.
C) Compilation of Results
The experimenters collected and aggregated the scores. Method 1 (M1) corresponds to GNNExplainer, and Method 2 (M2) corresponds to Logic. These results are reported in Table 5.3.